OpenStudy (anonymous):

If the area of a circle whose diameter is pi, is written as a times pi to the b power, what is the value of ab?

OpenStudy (anonymous):

no more drawings this time

OpenStudy (anonymous):

I got pi cubed/4

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so how could I rewrite this?

OpenStudy (anonymous):

rewrite what?

OpenStudy (anonymous):

pi cubed/4

OpenStudy (anonymous):

I need to rewrite as a times pi to the b power

OpenStudy (anonymous):

So, the area is π³/4. Rewrite it as π³/4 = aπ^b. Find the a and b.

OpenStudy (mr.math):

You know the area of a circle is \(\pi r^2\), where \(r\) is the radius of the circle. In this case we have \(r=\frac{\pi}{2}\). Thus the area, call it \(A\), is \(A=\pi(\frac{\pi}{2})^2=\frac{1}{4}\pi^3.\) Now if we set this \(A=a\pi^b=\frac{1}{4}\pi^3.\) Then we have \(a=\frac{1}{4}\) and \(b=3\). So \(ab=\frac{3}{4}\).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You asked me why a should be 1/4, not 4. It's because 4 is in denominator, so it is 1/4. π³/4 = (1/4)π³